There are some rabbits and some hutches. If eight rabbits are put in each hutch, one rabbit is left over. If eleven rabbits are put in each hutch, one hutch is left empty. How many rabbit hutches are there, and how many rabbits are there?
Answer by macston(5194)

.
H=number of hutches
.
8H+1=11(H-1)
8H+1=11H-11
12=3H
4=H
ANSWER 1: There are 4 hutches.
.
11(4-1)=11(3)=33
ANSWER 2: There are 33 rabbits.
.
CHECK:
8(4)+1=33
32+1=33
33=33
.